OpenStudy (mendicant_bias):

(ODE) Setting up a variation of parameters problem shortly. Having an issue somewhere down the line possibly with the Wronskian. Prompt posted below, along with workings so far.

OpenStudy (mendicant_bias):

\[y''-y=\sec(x).\]

OpenStudy (mendicant_bias):

\[\text{Characteristic Equation:} \ m^2-m=0; \ \ \ m(m-1)=0.\]

OpenStudy (mendicant_bias):

\[m=0, \ m=-1. \]

OpenStudy (mendicant_bias):

\[y_{h}=c_{1}e^0+c_{2}e^{-x}=c_{1}+c_{2}e^{-x}.\]

OpenStudy (mendicant_bias):

\[ W(e^{-x},1)= \begin{vmatrix} c^{-x}&1\\ -e^{-x}&0 \end{vmatrix} \]

OpenStudy (mendicant_bias):

\[W=-e^{-x} \neq 0,\]Thus, the solutions are linearly independent, at least. Now just to find them:

OpenStudy (mendicant_bias):

\[ W_1= \begin{vmatrix} 0&1\\ sec(x)&0 \end{vmatrix} =-sec(x).\]

OpenStudy (mendicant_bias):

\[ W(y_1,y_2)= \begin{vmatrix} e^{-x}&0\\ -e^{-x}&sec(x) \end{vmatrix} =e^{-x} sec (x)\]

OpenStudy (mendicant_bias):

Whoops, wasn't supposed to have that notation on the left but w/e.

OpenStudy (mendicant_bias):

@ganeshie8

OpenStudy (mendicant_bias):

\[\frac{W_{1}}{W}=u'(x)=\frac{-\sec(x)}{e^{-x}}=-e^{-x}\sec(x)\]

OpenStudy (mendicant_bias):

\[\frac{W_{2}}{W}=\frac{e^{-x}\sec(x)}{e^{-x}}=\sec(x)\]

ganeshie8 (ganeshie8):

heyy

OpenStudy (mendicant_bias):

Integrating both of those: \[\int\limits_{}^{}\sec(x)=\ln|(\sec(x)+\tan(x)|+C\]

OpenStudy (mendicant_bias):

Yoooo (lol)

OpenStudy (mendicant_bias):

What, did I make a mistake somewhere?

OpenStudy (mendicant_bias):

\[\int\limits_{}^{}e^{-x}\sec(x)= \text{(Integration by parts)} \dots\]

OpenStudy (mendicant_bias):

Says that the antiderivative can't be found...by an integral calculator.

OpenStudy (mendicant_bias):

@hartnn

OpenStudy (mendicant_bias):

OH, nevermind, beginning characteristic equation was wrong.

OpenStudy (mendicant_bias):

\[y''+y=\sec(x); \ \ \ m^2+1=0. \]

OpenStudy (mendicant_bias):

\[m = \pm \sqrt{-1}=\pm i\]